OpenStudy (anonymous):

Find the greatest 4 digits number which when divided by 18,24,30 and 36 leaves a remainder 17 (in each case).

OpenStudy (amoodarya):

a=18q+17 a=24k+17 a=30p+17 so 30|a-17 24|a-17 18|a-17 so [24,18,30]|a-17

OpenStudy (anonymous):

I need a way by which I can calculate answer to such question in 1-2 mins time. This type of questions appears in admission exam for MS(CS).

OpenStudy (anonymous):

no offense but @amoodarya I got no clue what you did back there. A little explanation please ?

OpenStudy (usukidoll):

that notation for the second one is 30 divides a -17 similarly 24 divides a-17 and 18 divides 1-17

ganeshie8 (ganeshie8):

Let \(a\) = required number Notice that \(a-17\) is divisible by each of the given integers, therefore \(a-17\) must also be divisible by their lcm, \(360\) : \[a-17 = 360k\]

OpenStudy (amoodarya):

[30,24,18]=[5.3.2, 2.2.2.3, 3.3.2]=2.2.2.3.3.5=\[2^3*3^2*5=360\\a-17=360k\\a=360k+17\\a_\min=360*3+17=1080+17\]

OpenStudy (anonymous):

@ganeshie8 and others thank you!

OpenStudy (usukidoll):

when did k =3 for the last line? just wondering. unless we have k = 3 because it's the lcm of 18,24,30, and 36.

OpenStudy (anonymous):

Lol I think @amoodarya assumed k = 3 :P

OpenStudy (usukidoll):

:S there is always a reason to let k = 3. isn't it?

OpenStudy (anonymous):

I can't think of any reasons to do so :| cause 3 isn't given in the question

OpenStudy (usukidoll):

the only thing I can think of is that 3 is the lcm of 18.24.30.36.. not to mention that's part of the 3 x table 3 x 6 =18 3 x 8 = 24 3 x 10 = 30 3 x 12 = 36 3 x 3 x 2 = 18 3 x 2 x 2 x 2 = 24 3 x 2 x 5 = 30 3 x 3 x 2 x 2 Each of them has at least 1 3

OpenStudy (anonymous):

yup thats true

OpenStudy (anonymous):

@amoodarya it'll be less confusing if you just type down a few words explaining the problem like ganeshie8 did. Thanks

OpenStudy (amoodarya):

\[a=360k+17 \ge 1000\\360k \ge 1000-17\\k \ge \frac{1000-17}{360}\\k \space is \space integer \\k \ge 3\]

OpenStudy (usukidoll):

where did you get the 1000 from?

OpenStudy (anonymous):

largest 4 digit number is 9999 Man Imma cry now seriously

OpenStudy (anonymous):

@amoodarya you're not helping sir rather you are making it more confusing and complex :|

OpenStudy (usukidoll):

yeah this is how my adviser would do... just be like ok this this this the end.

OpenStudy (anonymous):

LOL I hate such kind of teachers >:| Teacher's job is to explain the whole process....

OpenStudy (usukidoll):

yeah I got a B- for absolutely knowing 5.5 out of 8 chapters in Passage to Advanced Mathematics book

OpenStudy (amoodarya):

oh . sorry i had a little mistake i wrote it for smallest 4 digits so you do it for the greatest , like that

OpenStudy (anonymous):

@ganeshie8 can you make things simple its seriously frying my brain :|

OpenStudy (usukidoll):

wow I wish my retired professor was on here. He was awesome.

ganeshie8 (ganeshie8):

spoonfeeding is not good always, it would be more fun if you try and solve the problem on your own based on the given hints

OpenStudy (amoodarya):

\[a \leq 9999\\360k+17 \leq 9999\\k \leq \frac{9999-17}{360}\]

OpenStudy (usukidoll):

well what if the person made so many attempts and then got fed up?

ganeshie8 (ganeshie8):

where are the "many" attempts ?

OpenStudy (anonymous):

^ yes that's what I am doing for last 1 hour :|

OpenStudy (usukidoll):

k is less than or equal to 27.7222222222 something seems legit

OpenStudy (usukidoll):

"many" attempts? lol like using every theorem out there

ganeshie8 (ganeshie8):

teaching is not easy... there is no way for a teacher to know what the student has already tried right

OpenStudy (usukidoll):

showing off like what my professor did to me a while back doesn't help me understand either -_-

OpenStudy (anonymous):

^

ganeshie8 (ganeshie8):

try this : teach me how to multiply 3*5

OpenStudy (usukidoll):

ummm let's see. it's like writing 3 5 times and adding them up

OpenStudy (anonymous):

3*5 = 15 simple enough :O

ganeshie8 (ganeshie8):

how do you know i know addition ?

OpenStudy (usukidoll):

omg! take 3 rocks that's 1 batch then 3 rocks 3 rocks 3 rocks 3 rocks... add them all up

OpenStudy (anonymous):

alright I'll solve it myself .... no more arguments

ganeshie8 (ganeshie8):

where do i find rocks ? your explanation is confusing me

OpenStudy (usukidoll):

you find rocks outside!

OpenStudy (amoodarya):